a rectangle has an area of 900. the ratio of the l to w is 9 to 4. What is the length and width?
Answer by robertb(5830)

LW = 900.
L/W = 9/4. This equation gives L = 9W/4. This means (9W/4)*W = 900.
W^2 = 400,
W = 20, width is 20 units, and length is 45 units.